# SEMPO - Retrieving poles, residues and zeros in the complex frequency plane from an arbitrary spectral response [^1] This work was funded by the French National Research Agency ANR Project DILEMMA (ANR-20-CE09-0027)

### v1

@@ 3.2 Other forms of the singularity expansion
  • The Generalized Drude-Lorentz model (GDL) [44, 24] models the physical systems as sets of sub-units, each behaving as oscillators described by classical mechanics. The sub-units can be identified as the microscopic-scale particles constituting the system.
Each of these forms is truncated and written in a way that is compatible with the optimization problem, similarly to the classical SEM in Equation (26).
@@
where we assume that there is a pole 00 at the origin, associated with a residue $\mathrm{i}r_{0}$. Let us stress that the absolute value ”|.||.|| . |” is used to impose the stability of all the poles, but it can be removed to allow for unstable effective poles. The following expression of the GDL is obtained:
@@
We apply it to the case of the lattice of gold nanodisks presented in Figure 5 (a). The studied signal is the transmission coefficient $h(\omega)=\tilde{t}(\omega)$ at normal incidence. The reconstruction shown in Figure 5 (b,c) is obtained with $M_{I}=1$ imaginary pole, $M_{C}=8$ complex pairs of poles and $\mathbf{\alpha}=(1,0,0,0)$. The weight coefficients are such that more importance is given to the real part $h(\omega)$ which shows more complex variations than the imaginary part. We obtain a relative $L_{2}$ error of $6.31\times 10^{-4}$, over a spectral range extending from $100$nm to $900$nm.
Figure 5: (a) Lattice of gold nanodisks with height $h=50$nm, diameter $d=80$nm and period $p=500$ nm, bearing on a substrate of glass. The signal is the $0^{th}$-order transmission coefficient $h(\omega)=t_{00}(\omega)$ for a normal incidence in air. (b) Reconstruction of $h(\omega)$ with the GDL in Equation (43) (red curve), and compared to simulated data (blue markers). A relative $L_{2}$ error $e^{(2)}=6.31\times 10^{-3}$ is obtained. (c) Distribution of the poles in the complex $\omega$ plane.
@@ 4 Combined approach
Figure 6: Influence of the frequency sampling in the ADC method on the relative $L_{2}$ error. The signal is the reflection coefficient of a slab of $\mathrm{TiO}_{2}$ $h(\omega)=\tilde{r}(\omega)$ of thickness $d=260$ nm, for a TM illumination at $17\degree$ coming from the air superstrate. $N$ frequencies are uniformly picked between $0.15$ and $15.6\times 10^{15}$ rad/s. (a) Reconstruction of $h(\omega)$ using $N=40$ (blue), $N=60$ (orange) and $N=80$ (green) frequencies, and compared to the target (black). (b) Evolution of the relative error with the number of frequencies, with $N$ going $5$ by $5$ from $20$ to $100$.
The ADC method is faster than the auto-differentiation approach. However, it can be difficult to maintain a high accuracy while opting for a Hermitian-symmetric and stable solution, as previously shown. In addition, the method does not work well in large spectral windows showing rapid variations of the signals, and is sensitive to the sampling of data. While the latter is addressed with what is known as the adaptive Cauchy method [16], the former is not easily solved for. We illustrate both problems in Figure 6, where we apply the ADC method to retrieve the poles and zeros of the reflection coefficient of a slab of $\mathrm{TiO}_{2}$ $h(\omega)=\tilde{r}(\omega)$, of thickness $d=260$nm, with a TM illumination at $17\degree$ in air. $N$ frequencies are uniformly picked between $0.15$ and $15.6\times 10^{15}$ rad/s. We observe in Figure 6 (a) that fast dynamics such as the one around $5\times 10^{15}$ rad/s are difficult to capture, whatever the value of $N$. In Figure 6 (b) shows that the relative $L_{2}$ does not systematically lower when $N$ is increased.
@@
Figure 8: (a) Reflection coefficient $h(\omega)=\tilde{r}(\omega)$ of a slab of $\mathrm{TiO}_{2}$ of thickness $d=260$ nm, illuminated from the air superstrate with an incidence of $17\degree$, in the TM polarization. It is reconstructed using either the SEM (blue curve), or the combined ADC and SEM approach (red curve), and compared to simulated data (black markers). A relative $L_{2}$ error $e^{(2)}=5.73\times 10^{-2}$ is obtained with the SEM approach, and $e^{(2)}=1.08\times 10^{-2}$ for the combined approach. The latter thus performs better, and is obtained in half of the time needed for the former. (b) Distribution of the poles in the complex $\omega$ plane, for the SEM (blue markers) and the combined approach (red markers). Fewer singularities are required with the ADC+SEM approach to achieve a better reconstruction than with the SEM.
@@
In every sub-window, the poles with a weight $q_{w,\ell}$ below an arbitrary threshold $q_{0}=68$ are filtered out. We then concatenate all the remaining poles and residues into unique distributions used as initial parameters in the auto-differentiation approach.
@@
We compare this combined approach to the classical SEM in Figure 8, for the same reflection coefficient as in Figures 6 and 7, i.e. the case of a slab of $\mathrm{TiO}_{2}$ In Figure 8 (a), we observe a slightly more accurate reconstruction with the combined ADC+SEM approach than with the SEM only. Let us stress that the optimization process takes, in this case, half as long with the combined approach than with just the SEM. Another significant advantage is that fewer singularities are needed with the ADC+SEM approach than with the SEM only, as shown in Figure 8 (b), where the singularities associated to both approaches are shown in a frequency window of the complex $\omega$ plane.

### v2

@@ 3.2 Other forms of the singularity expansion
  • The Generalized Drude-Lorentz model (GDL) [47] models the physical systems as sets of sub-units, each behaving as oscillators described by classical mechanics. The sub-units can be identified as the microscopic-scale particles constituting the system.
@@
where we assume that there is a pole 0 at the origin, associated with a residue $\mathrm{i}r_{0}$.
Figure 5: (a) Lattice of gold nanodisks with height h=50h=50nm, diameter d=80d=80nm and period $p=500$ nm, bearing on a substrate of glass. The signal is the $0^{th}$-order transmission coefficient $h(\omega)=t_{00}(\omega)$ for a normal incidence in air. (b) Reconstruction of $h(\omega)$ with the GDL in Equation (40) (red curve), and compared to simulated data (blue markers). A relative $L_{2}$ error $e^{(2)}=6.31\times 10^{-3}$ % is obtained. (c) Distribution of the poles in the complex $\omega$ plane.
Let us stress that the absolute value ”|.||.|” is used to impose the stability of all the poles, but it can be removed to allow for unstable effective poles. The following expression of the GDL is obtained:
@@
We apply it to the case of the lattice of gold nanodisks presented in Figure 5 (a). The studied signal is the transmission coefficient $h(\omega)=\tilde{t}(\omega)$ at normal incidence. The reconstruction shown in Figure 5 (b,c) is obtained with $M_{I}=1$ imaginary pole, $M_{C}=8$ complex pairs of poles and $\mathbf{\alpha}=(1,0,0,0)$. The weight coefficients are such that more importance is given to the real part $h(\omega)$ which shows more complex variations than the imaginary part. We obtain a relative $L_{2}$ error of $6.31\times 10^{-4}$ %, over a spectral range extending from 100100nm to 900900nm.
@@ 4 Combined approach
Figure 6: Influence of the frequency sampling in the ADC method on the relative $L_{2}$ error. The signal is the reflection coefficient of a slab of $\mathrm{TiO}_{2}$ $h(\omega)=\tilde{r}(\omega)$ of thickness $d=260$ nm, for a TM illumination at $17\degree$ coming from the air superstrate. NN frequencies are uniformly picked between $0.15$ and $15.6\times 10^{15}$ rad/s. (a) Reconstruction of $h(\omega)$ using $N=40$ (blue), $N=60$ (orange) and $N=80$ (green) frequencies, and compared to the target (black). (b) Evolution of the relative error with the number of frequencies, with NN going 55 by 55 from $20$ to $100$.
The ADC method is faster than the auto-differentiation approach. However, it can be difficult to maintain a high accuracy while opting for a Hermitian-symmetric and stable solution, as previously shown. In addition, the method does not work well in large spectral windows showing rapid variations of the signals, and is sensitive to the sampling of data. While the latter is addressed with what is known as the adaptive Cauchy method [39], the former is not easily solved for. We illustrate both problems in Figure 6, where we apply the ADC method to retrieve the poles and zeros of the reflection coefficient of a slab of $\mathrm{TiO}_{2}$ $h(\omega)=\tilde{r}(\omega)$, of thickness d=260d=260nm, with a TM illumination at $17\degree$ in air. $N$ frequencies are uniformly picked between $0.15$ and $15.6\times 10^{15}$ rad/s. We observe in Figure 6 (a) that fast dynamics such as the one around $5\times 10^{15}$ rad/s are difficult to capture, whatever the value of $N$. In Figure 6 (b) shows that the relative $L_{2}$ does not systematically lower when $N$ is increased.
@@
Figure 8: (a) Reflection coefficient $h(\omega)=\tilde{r}(\omega)$ of a slab of $\mathrm{TiO}_{2}$ of thickness $d=260$ nm, illuminated from the air superstrate with an incidence of $17\degree$, in the TM polarization. It is reconstructed using either the SEM (blue curve), or the combined ADC and SEM approach (red curve), and compared to simulated data (black markers). A relative $L_{2}$ error $e^{(2)}=5.73\times 10^{-2}$ % is obtained with the SEM approach, and $e^{(2)}=1.08\times 10^{-2}$ % for the combined approach. The latter thus performs better, and is obtained in half of the time needed for the former. (b) Distribution of the poles in the complex $\omega$ plane, for the SEM (blue markers) and the combined approach (red markers). Fewer singularities are required with the ADC+SEM approach to achieve a better reconstruction than with the SEM.
@@
In every sub-window, the poles with a weight $q_{w,\ell}$ below an arbitrary threshold $q_{0}=68$ % are filtered out. We then concatenate all the remaining poles and residues into unique distributions used as initial parameters in the auto-differentiation approach.
@@
We compare this combined approach to the classical SEM in Figure 8, for the same reflection coefficient as in Figures 6 and 7, i.e. the case of a slab of $\mathrm{TiO}_{2}$ . In Figure 8 (a), we observe a slightly more accurate reconstruction with the combined ADC+SEM approach than with the SEM only. Let us stress that the optimization process takes, in this case, half as long with the combined approach than with just the SEM. Another significant advantage is that fewer singularities are needed with the ADC+SEM approach than with the SEM only, as shown in Figure 8 (b), where the singularities associated to both approaches are shown in a frequency window of the complex $\omega$ plane.
5 Pole retrieval efficiency
5.1 Performance evaluation criteria
The singularity expansion in Equation (1) provides an exact expression of any physical signal $h(\omega)$ in terms of the potentially infinite sets of complex poles $p^{(\ell)}$ and associated residues $r^{(\ell)}$. If we assume that $h(\omega)$ is fully known on an open set $\Omega=\left]\omega_{0},\omega_{1}\right[$ of $\mathbb{R}$, then the identity theorem states that there is a unique singularity expansion that matches $h(\omega)$ on $\Omega$, effectively providing an expansion of $h(\omega)$ into the complex frequency plane and ensuring the unicity of the poles, residues and non-resonant term. If we are able to reduce the error between a singularity-expanded function and $h(\Omega)$ to zero, we then obtain the singularity expansion of $h(\omega)$, which we directly denote as $h(\omega)$.
Three problems arise in practice.
  1. (i)
    Any prior information is obtained by measuring or simulating $h(\omega)$ over a discrete set of frequencies, i.e. the aforementioned $N$ frequencies $\omega_{n}$. The identity thus no longer holds. Nevertheless, when the spacing between subsequent frequencies is small enough compared to the spectral range of interest, we implicitly assume that this first problem is negligible and that we can still use the identity theorem.
  2. (ii)
    Although the definition of an error function, such as the relative $L_{2}$ error, is possible, it is numerically impossible to construct a singularity expanded function with an infinite number of poles. This quickly leads to a truncation such as the one considered in Equation (3) for the SZF rather than the SEM expression. Consequently, the error cannot be reduced to 0, and the unicity cannot be guaranteed. Adding more singularities does not result in a convergence towards the actual function, even though a better approximation in terms of error reduction.
  3. (iii)
    Measurements or simulations are always associated with errors, even though they can be extremely small compared to the actual value in the case of simulations. The result is that the poles (and zeros or residues) that we attempt to retrieve are those of a close approximation of the actual function rather than this function $h(\omega)$ itself.
Figure 9: (a) Reconstruction of the test function $h(\omega)$ (cyan curve) described in Table 1 using the ADC method (black dotted curve) on a sample of $30$ points (black crosses), without imposing the Hermitian symmetry, with a stability constraint coefficient $q_{0}=0.05$, with a maximum of $20$ poles, and with a maximum difference of $4$ between the numbers of poles and zeros. The poles of the target function (Original poles, blue) and those retrieved by the Cauchy method (Retrieved poles, red) are shown in the log-amplitude map, in the complex $\omega$ plane. (b) A pole $p=(9.001-.71\mathrm{i})\times 10^{15}$ rad/s is added at high frequencies, close to an original pole of $h(\omega)$. This results in an averaged retrieved pole rather than the retrieval of both poles. (c) A pole $p=(2.421-.001\mathrm{i})\times 10^{15}$ rad/s is this time added at low frequencies, resulting in a similar effect, i.e. one average retrieved poles for two original poles. (d) A Gaussian noise is added to the sampled data, maintaining a Signal-to-Noise Ratio (SNR) of $30$. Only one pole, associated with the peak at around $2\times 10^{15}$rad/s, is accurately retrieved.
As we attempt to retrieve the poles and zeros or poles and residues, we are only able to recover an approximation of the function over a finite spectral window, with no guarantee that the obtained parameters offer a valid expansion of the data into the complex frequency plane. This is illustrated in Figure 9, where we apply the ADC method in four different situations on a test function $h(\omega)$ defined as
$h(\omega)=\sum_{\ell=1}^{5}\frac{r_{\ell}}{\omega-p_{\ell}},$ (47)
with the residues $r_{\ell}$ and poles $p_{\ell}$ given in Table 1.
Test function $h(\omega)$
Index $\ell$ Pole ($\times 10^{15}$rad/s) Residue ($\times 10^{15}$rad/s)
11 $2-2\mathrm{i}$ $e^{-\frac{1}{9}\mathrm{i}\pi}$
22 $2.2-2.3\mathrm{i}$ $e^{\frac{1}{9}\mathrm{i}\pi}$
33 $2.42-0.002\mathrm{i}$ $e^{\frac{17}{180}\mathrm{i}\pi}$
44 $5-2\mathrm{i}$ $e^{\frac{1}{9}\mathrm{i}\pi}$
55 $9-0.7\mathrm{i}$ $e^{\frac{1}{6}\mathrm{i}\pi}$
Table 1: Complex poles and residues of the test function $h(\omega)$ used in Equation (47).
In Figure 9 (a), we consider the exact function and show that the target poles and residues are retrieved. In Figure 9 (b) we add a pole close to the one outside of the sampled data, making it difficult to discriminate between the two. This results in an averaging of the poles (but not the residues). In Figure 9 (c), a similar effect is obtained when adding a pole close to the one with the smallest imaginary part to $h(\omega)$. In Figure 9 (d) we randomly add noise to the data, with a noise-to-signal ratio of less than $1\%$ for each value $h(\omega_{i})$, resulting in a combination of the two previous effects.
The actual poles hold important and rich information regarding a physical signal since they strongly contribute to its resonances [31]. It is therefore necessary to be able to discriminate, amongst all the retrieved poles, between the natural poles which would be obtained if the identity theorem could hold, i.e. with infinite precision over a continuous set of frequencies, and the effective poles, which result from the introduction of errors and only help in reducing the error between the truncated singularity expansion or SZF and the measured or simulated data.
Under these considerations, we evaluate the efficiency of algorithms through four scores:
  • The overall efficiency estimated via the relative $L_{2}$ precision $p_{r}^{(2)}=1-e^{(2)}$.
  • The physical relevancy of the solutions through the ratio of Hermitian-symmetric to total poles $\rho_{\mathrm{herm}}=\frac{\mathrm{Hermitian~poles}}{\mathrm{retrieved~poles}}$.
  • The stability of the solution with the ratio of stable to total poles $\rho_{\mathrm{stab}}=\frac{\mathrm{stable~poles}}{\mathrm{retrieved~poles}}$.
  • The ratio of natural to total poles $\rho_{\mathrm{nat}}=\frac{\mathrm{natural~poles}}{\mathrm{retrieved~poles}}$.
The computation of the first three scores is straightforward given a target curve and target poles. Nonetheless, the identification of natural poles is a complex problem. We calculate $\rho_{\mathrm{nat}}$ in two steps. Given a set of target natural poles $\mathcal{P}=\{p_{1},...,p_{M_{p}}\}$ and a set of retrieved poles $\mathcal{Q}=\{q_{1},...,q_{M_{q}}\}$, we calculate a relative distance matrix $\mathbf{D}=[D_{ij}]$ with $1\leq i\leq M_{p}$ and $1\leq j\leq M_{q}$, defined as
$D_{ij}=\min\left(\frac{|p_{i}-q_{j}|}{|p_{i}|},\frac{|p_{i}-q_{j}|}{|q_{i}|}\right)$ (48)
Using $\mathbf{D}$, we identify all the pairs of poles $(p_{i},q_{j})$ separated by a relative distance $\delta_{ij}$ shorter than an arbitrarily chosen threshold $\delta_{0}=10\%$. If $p_{i}$ and $q_{j}$ are close enough, we then decide on whether they are identical based on similarity of the resonances they would bring about in a response function if they were isolated. To do so, we first calculate the quality functions $\eta_{i}$ of $p_{i}$ and $\eta_{j}$ of $q_{j}$, which can be interpreted here as similar as the quality factors of the isolated resonances, but calculated over a spectral range [31]:
$\eta_{i}(\omega)=\frac{\mathrm{i}}{4}\left[\frac{\omega}{\omega-p_{i}}-\frac{\omega}{\omega-\overline{p_{i}}}\right]$ (49)
$\eta_{j}(\omega)=\frac{\mathrm{i}}{4}\left[\frac{\omega}{\omega-q_{j}}-\frac{\omega}{\omega-\overline{q_{j}}}\right]$ (50)
We then compute the standard deviation $\sigma_{ij}$ of the absolute difference of the quality functions evaluated at the $N$ sample frequencies $\omega_{n}$:
$\sigma_{ij}=\mathrm{st.dev.}\left(\{\left|\eta_{i}(\omega_{n})-\eta_{j}(\omega_{n})\right|\}_{1\leq n\leq N}\right)$ (51)
The value of $\sigma_{ij}$ can be understood as a quantification of the maximum influence of the difference of the resonances brought about by the isolated poles $p_{i}$ and $q_{j}$. If $\sigma_{ij}<2$, i.e. if the difference of the resonances has a quality factor smaller than $2$, then we assume that $p_{i}$ and $q_{j}$ are identical, in which case the retrieved pole $q_{j}$ is counted as a natural pole if $p_{i}$ is a known target pole.
5.2 Algorithm benchmarking
Using these criteria, we evaluate the performances of the ADC method and the auto-differentiation (AutoDiff) approach of SEMPO, which we compare to the original Cauchy method [58], as well as of two state-of-the-art methods, i.e. the AAA algorithm [55, 56] and the Vector Fitting (VF) approach [53, 72, 73]. The test function we want to fit is the Hermitian-symmetric version of function introduced in Equation (47):
$h(\omega)=\sum_{\ell=1}^{5}\left[\frac{r_{\ell}}{\omega-p_{\ell}}-\frac{\overline{r_{\ell}}}{\omega+\overline{p_{\ell}}}\right],$ (52)
where the residues $r_{\ell}$ and poles $p_{\ell}$ are those given in Table 1. We uniformly sample $h(\omega)$ to obtain a fitting dataset of $35$ frequencies $\omega_{i}$ ranging from $7\times 10^{15}$rad/s to $1\times 10^{15}$rad/s and their associated values $h_{i}=h(\omega_{i})$. While this sample is used to run the algorithms, the relative $L_{2}$ precision is then calculated for a sample of $100$ points in the same spectral window. We test the performances of the different algorithms on noisy data $\tilde{h}_{i}$, obtained by adding Gaussian noise to $h_{i}$, with an SNR ranging from $10\log_{10}(50)$ to $10\log_{10}(1000)$. The noisy data read as
$\tilde{h}_{i}=h_{i}+\sigma(b_{R,i}+\mathrm{i}b_{I,i}),$ (53)
where
$b_{R,i},b_{I,i}\sim\mathcal{N}(0,1).$ (54)
We write the SNR as
$SNR=10\log_{10}\left(\frac{\displaystyle\sum_{i}|{h_{i}}|^{2}}{\displaystyle\sigma^{2}\sum_{i}|{b_{R,i}+\mathrm{i}b_{I,i}}|^{2}}\right)$ (55)
allowing us to set the value of $\sigma$ via
$\sigma=10^{-SNR/20}\sqrt{\frac{\sum_{i}|{h_{i}}|^{2}}{\sum_{i}|{b_{R,i}+\mathrm{i}b_{I,i}}|^{2}}}.$ (56)
The SNR is set successively to $10\log_{10}(50)$, $10\log_{10}(100)$ and $10\log_{10}(1000)$ during the benchmarking.
We perform a parameter sweep for all the aforementioned pole retrieving methods. Each set of parameter is tested with different levels of Gaussian noise in the test data as the SNR is increased from $10\log_{10}(50)$ to $10\log_{10}(1000)$. The VF approach works by fitting the data with an expression similar to the truncated singularity expansion in Equation (1), to which is optionally added a term proportional to the frequency $\omega$. We run the algorithm without this additional term, and only sweep through the number of complex Hermitian-symmetric poles and the number of purely imaginary poles:
  • The number of complex Hermitian-symmetric poles ranges from $4$ to $18$ with a step of $2$.
  • The number of purely imaginary poles ranges from 0 to $4$ with a step of $1$.
The AAA algorithm fits the data with a stable rational approximation that takes the form of a ratio of partial fractions. This rational approximation can be recast as a truncated SZF, allowing the AAA algorithm to provide the poles and zeros of the fitted function. Only one parameter corresponding to the maximum number of terms in the partial fractions is swept through:
  • The maximum number of terms ranges from $2$ to $14$ with a step of $1$.
In our implementation of the original Cauchy method described in Section 2.1, only the maximum number of poles is swept through:
  • The maximum number of poles ranges from $8$ to $20$ with a step of $2$.
We run the ADC method with an imposed Hermitian symmetry, and a maximum difference of $4$ between the numbers of poles and zeros. We sweep through the maximum number of poles and the stability constraint coefficient $q_{0}$:
  • The maximum number of poles ranges from $8$ to $20$ with a step of $2$.
  • The stability coefficient $q_{0}$ takes $10$ evenly spaced values from 0 to $0.05$.
For the auto-differentiation approach, the hyperparameters are fixed: the learning rate is set to .007.007, and the number of iterations to $22000$, which is about two third of the value classically used in the optical structures of the previous structures. The swept-through parameters are the number of complex Hermitian-symmetric poles and the number of purely imaginary poles:
  • The number of complex Hermitian-symmetric poles ranges from $2$ to $10$ with a step of $1$.
  • The number of purely imaginary poles goes from 0 to $2$ with a step of $1$.
Let us point out that we do not consider here the combined chained-ADC and auto-differentiation approach as it is primarily used in wide spectral range exhibiting rapid variations. The chained ADC method presents no advantage with the test function used for the benchmarking as it introduces several singularities which cannot be all removed at the seams of the sub-windows, despite the merging routine described in Section 4.
Other methods operating in the temporal domain such as the harmonic inversion [52] or the Prony-SS decomposition method [74] are not tested here, since they involve a Fourier transform to obtain temporal data, which might lead to additional sources of error difficult to account for. Nonetheless, these temporal approaches work by fitting time-decaying exponentials corresponding to the analytical inverse Fourier transform of the singularity expansion. We thus expect them to perform similarly to SEMPO. In addition, the benchmarking scripts are available on the Zenodo repository to allow for additional performance evaluations with other test functions or algorithms.
Figure 10: Benchmarking results of the AAA algorithm, VF algorithm, original Cauchy method (CAUCHY), ADC method and auto-differentiation (AutoDiff) approach, for an SNR of $10\log_{10}(50)$, $10\log_{10}(100)$ or $10\log_{10}(1000)$. (a) Average relative $L_{2}$ precision. (b) Average Hermitian-symmetric pole ratio. (c) Average stable pole ratio. (d) Average natural pole ratio. The averaging is performed over the set of parameters for each SNR level.
As previously mentioned, we filter out all the algorithm runs performed with an SNR $\leq 10\log_{10}(10)$, and calculate the four scores defined in the previous Subsection 5.1 for each of the remaining runs. The results are presented in Figure 10 where they are grouped up based on the SNR and the algorithm or method used. The average relative $L_{2}$ precision $p_{r}^{(2)}$ is in Figure 10 (a), the Hermitian ratio $\rho_{\mathrm{Herm}}$ in Figure 10 (b), the stability ratio $\rho_{\mathrm{stab}}$ in Figure 10 (c), and the Natural pole ratio $\rho_{\mathrm{nat}}$ in Figure 10 (d). The averaging is performed over the set of parameters for each SNR level. For a given SNR, we generate $50$ datasets by drawing $50$ times the noise. The scores for a given algorithm, with a given set of parameters is thus an average over these $50$ drawings. Let us point out that the stability ratio and Hermitian ratio are biased towards SEMPO since it has been developped in order to account for these physical constraints. Their presence is however helpful in understanding the advantage of some methods compared to others.
In terms of average $L_{2}$ precision, we find that the VF algorithm, the ADC method and the AutoDiff approach perform similarly, as observed in Figure 10 (a). The ADC ranks first at SNR levels below $10\log_{10}(1000)$, but is equivalent to the AutoDiff and VF as the SNR reaches $10\log_{10}(1000)$. This result is in agreement with the design of the ADC method, which builds upon the classical Cauchy method while minimizing the relative error. Let us point out that by increasing the value of the hyperparameter corresponding to the number of iterations from $20000$ (used for the benchmarking) to $35000$ (usual parameter), the average precision of the AutoDiff can be increased at all SNR. This shows the sensitivity of the AutoDiff approach to the hyperparameters.
The high precisions of the VF algorithm and ADC methods comes at the expense of non-physical solutions. The VF algorithm does not comply with the Hermitian-symmetry as shown in Figure 10 (b), while the stability of the solutions is not fully ensured in the ADC method as seen in Figure 10 (c). Only the AutoDiff approach yields solutions fully respecting both criteria.
Finally, we observe that the AutoDiff approach offers a more robust way of retrieving the actual poles of a system as indicated by the higher natural pole ratio in Figure 10 (d) with respect to other methods, at all SNR. This ratio is obtained by dividing the number of natural poles by the number of retrieved poles instead of actual poles. This allows us to reward methods which provide fewer poles, making it more likely to identify them when the target is unknown, i.e. in classical applications.
Overall, the ADC method provides a good alternative to the VF algorithm for obtaining a Hermitian-symmetric solution with high precision, but it lags behind in stability. On the other hand, the AutoDiff approach provides both accurate solutions and Hermitian-symmetric, stable poles, while showing the best performance in the retrieval of natural poles of the system bringing about the signal.